OpenStudy (anonymous):

14 474 51415 1235321 1111?1111. find the missing number ? a.10 b.1 c.2 d.4

OpenStudy (anonymous):

a

OpenStudy (anonymous):

no......its given option b.1 in the answer sheet

OpenStudy (anonymous):

but couldnt solve

OpenStudy (anonymous):

I still say a. 10

OpenStudy (anonymous):

howz that?

OpenStudy (anonymous):

can u explain plz?

OpenStudy (anonymous):

look for a pattern in the rows

OpenStudy (anonymous):

i think b

OpenStudy (anonymous):

i dont think so ;

OpenStudy (anonymous):

ya its b

OpenStudy (anonymous):

but how did u do that

OpenStudy (anonymous):

is the top row 1, 4 or 14?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

followed by 474

OpenStudy (anonymous):

well, if top row is fourteen, 2nd row is four-seven-four, etc, then I say a. 10 :)

OpenStudy (anonymous):

still didnt get ur point./

OpenStudy (anonymous):

add the numbers. 14=14 4+7+4=15 5+1+4+1+5=16 etc

OpenStudy (anonymous):

does this patterrn follow all the row?

OpenStudy (anonymous):

i mean if so than 1+4=5 should be in st row

OpenStudy (anonymous):

isnt it ?

OpenStudy (anonymous):

thats why i asked if it was one-four or fourteen

OpenStudy (anonymous):

it can be taken either in 1 4 or 14.....but all should be in same pattern/

OpenStudy (anonymous):

as i can't see the paper or wherever you are getting this from, i dunno

OpenStudy (anonymous):

well this was the question asked to me in the medical entrance test

OpenStudy (anonymous):

anyway i respect ur logic above...............but it was not the one that suits this question...

OpenStudy (anonymous):

sorry.

OpenStudy (anonymous):

i mean that was not the logic the question want...........i guess.:)

OpenStudy (anonymous):

but i guess the question marks should be given in last position of last row as i guess.